### v1

@@ 5 Discussion
The JWST observations reveal a detailed hourglass-shaped bipolar outflow in the MIR broadband images (see Fig. 1). The images exhibit multiple nebulous features consisting of warm and hot layers of gas and dust generated within the outflow cavity. Our spectral maps reveal that the emission includes molecular, ionized and scattered light. In spite of such detailed composition, we do not see any trace of a large scale collimated jet in the broadband images. Such an absence of a large jet signature is distinctive and unique for L1527 given its evolutionary stage. In contrast, recent JWST observations of several other Class 0/I protostars like HH 211, HH 46, TMC 1, IRAS 16253-2429, have shown to have prominent jet emission. Their bulk of the emission is particularly traced in high-J H2 and [Fe ii] lines (Ray et al., 2023; Nisini et al., 2024; Tychoniec et al., 2024; Narang et al., 2024). We see that most of L1527’s dominant emission comes from the wide angle outflow across both molecular and ionized emission. However, the detection of double peaked ionized lines with the HV component having velocities 100 km s-1gives compelling evidence for the presence of a jet. This jet signature spans less than 500 au from the protostar and is fainter in relation to the jets from other sources. In addition, when compared to the large scale outflow structures (3000 au), the jet is small and covers only a fraction of the region, thereby concealing its visibility in the broadband images.
@@
Double peaked emission lines with kinematics similar to L1527 have been observed in other protostellar outflows. For instance, HH 34 presents a collimated jet showing two velocity components both in molecular and ionized lines (Nisini et al., 2016). Using VLT observations of H2 and [Fe ii] lines, Garcia Lopez et al. (2008) showed that HH 34 LV component is distributed at 15-50 km s-1surrounding a fast axial jet (90-140 km s-1). Likewise, in DG Tau, [Fe ii] and [Ne iii] double peaked lines were observed with VLT: SINFONI and X-shooter showing a similar velocity separation (Agra-Amboage et al., 2011; Liu et al., 2016). Many other sources exhibit the same pattern, for e.g. SVS 13, HH 211, HH 46 (Takami et al., 2006; Tabone et al., 2017; Garcia Lopez et al., 2010). Such profiles are rarely seen in Class 0 protostars due to their dense envelopes, but more notable in Class II protostars due their moderate extinction. A unifying scenario in all these observations is that the LV component is associated to a slower wider wind, while the HV component traces the faster jet, thereby pointing to a stratified structure in the outflow.
@@
The observed double peak lines can also be interpreted within the framework of shock models. The coexistence of the LV and HV components in the ionized lines suggests a stratified velocity structure within predominantly -type shock environment. Our comparison to models shows that moderate velocity ( km s-1) -type shocks illuminated by non-zero UV field best reproduce the observed emission. This suggests that the bulk of the line excitation is driven by compact, dissociative shock fronts likely associated to the disk wind. The high [Ne ii]/[Ne iii] ratio across most of the emitting region indicates soft ionization overall, consistent with moderate-velocity shocks where UV photons generated in the shock front contribute to the ionization. However, the clear detection of HV component in [Ar ii], [Ne iii] and [Fe ii] specifically indicates a localized contribution from a faster jet. The HV component, particularly in [Ne iii] is significant as ionization of Neon requires energies 40 eV, typically only achieved via fast dissociative -type shocks (80 km s-1) (Hollenbach & McKee , 1989; Shang et al., 2010), as strong X-ray/UV photoionization is unlikely in L1527. The inferred jet velocity (100 km s-1) in L1527 is consistent with predictions of a shock-excited jet emerging from the inner disk, suggesting the existence of a high-velocity jet core. All together, these observations support a stratified jet-outflow structure in which the fast jet core produces HV component in the ionized lines, while the moderate velocity -type shock with intermediate UV field contributes to the broader LV component throughout the outflow.
@@
Radio observations have been used to detect jets based on the continuum emission arising from the thermal Bremsstrahlung radiation (free-free emission) (e.g. Anglada et al., 1998; Furuya et al., 2003). An earlier study of L1527 by Reipurth et al. (2004) using VLA ’A’ configuration at 3.6 cm observed an extended continuum jet in the eastern region close to the protostar. The jet emission was extended to about 150 au and was seen to curve towards the south on its leading end (see black contours in Fig. 11). This detection was not taken into consideration, as it was speculated that it may be due to a binary as proposed by Loinard et al. (2002). However, based on recent high-sensitivity observations it was shown that L1527 is not a binary (e.g. Nakatani et al., 2020; Sheehan et al., 2022). Correlating the morphology of the radio jet to the MIRI jet, we find similarity in the distribution. The jet axis matches spatially extending towards the east and bending/curving towards the south. The curving of the jet could be due to wiggling/precession or an interaction with a dense layer of gas. A key feature that can be interpreted from both observations is the proper motion of the jet. The radio jet extends approximately to about 150 au, whereas the MIRI jet emission extends up to 500 au. There is a difference of about 350 au in length between both the emissions which have a time difference of 20 years. Based on this, a rough estimate of the average tangential velocity of the jet would lead to around km s-1. A main caveat in this estimate is that both data sets are observed at different wavelengths, therefore the difference in the jet length may also be associated to the emission from different layers of gas.
@@
More recently, van’t Hoff et al. (2023) carried out ALMA observations and detected a noticeable SiO emission on the western side (see blue contours in Fig. 11). SiO predominantly traces high density ejection material that are associated to jets in the innermost regions of Class 0/I protostars (e.g. HH 212; Lee , 2020). The SiO emission observed in L1527 is compact and positioned very close to the protostar (10-50 au). Most importantly, it is blue shifted and exhibits higher velocity when compared to ALMA CO observations. When considered alongside our MIRI jet emission on the eastern side, this detection strongly supports the bipolar nature of jet in L1527. In addition their velocities and position confirm that the eastern region is red shifted while the western region is blue shifted consistent with the [Ne ii] velocity maps (see Fig. 6). van’t Hoff et al. (2023) also examined a large scale 12CO map and found hints of a collimated CO emission on the western side at a distance of around 500-800 au. This emission was seen to be more resolved, flowing along the outflow axis, but had lower velocity than SiO emission. Therefore, it could not be confirmed as a large scale jet.
@@
While the L1527 outflow exhibits a clear bipolar structure, the jet emission traced by HV component and SiO appear asymmetric and poorly distributed. The confinement of the HV component to one region may reflect the effects due to combination of inclination, asymmetries in the medium, or an episodic accretion. Asymmetric jets have been observed in other deeply embedded sources such as L1448-C (Dionatos et al., 2009), BHR 71 (Gusdorf et al., 2015), HH 46/47 (Zhang et al., 2016) and TMC 1 (Tychoniec et al., 2024), where external density gradients or variable accretion lead to predominantly one-sided jet activity. The inhomogeneity of the medium significantly changes the excitation conditions leading to emission in only certain tracers for each region. Many theoretical simulations lend strong support to this interpretation. For instance, Raga et al. (2002) modeled jets in young stars and showed how early-stage jets can produce compact, asymmetric jet features, particularly when interacting with clumpy or inhomogeneous envelopes. Hansen et al. (2016) conducted 3D MHD simulations of Class 0 jets and found that early jet interactions with over-dense clumps can produce complex asymmetric shock structures. In addition, such interactions also produce number of reverse shocks along the flow, which seems to be the case in L1527 (see Fig. 1). Moreover, it is believed that magnetic fields play a crucial role in shaping the jet structures (Cerqueira & Gouveia Dal Pino , 2001; Pudritz & Ray , 2019). Variations in magnetic field topology can lead to jets with uneven or asymmetric outflows. Different configurations of magnetic fields, such as misalignments between the field and the rotation axis due to warped disk, or non-uniform field structures, can result in one side of the jet being less collimated or more powerful than the other.
@@
Another factor attributing to L1527 jets morphology is the evolutionary stage. While L1527 is classified as a Class 0/I transitional object, its jet appears unusually weak and compact. This indicates that the jet activity may have declined or it undergoes episodes of periodic mass-loss. Such temporal evolution has been seen in sources like L1448-C, IRAS 16293-2422, Serpens South C7, where compact jets coexist with broader molecular flows, and signs of episodic accretion are common (Bachiller et al., 1990; Kristensen et al., 2011; Plunkett et al., 2015). Spitzer IRAC observations of L1527 taken 12 months apart show variability in total flux and variability in the emission in each outflow lobes (Tobin et al., 2008). Such variations are typically associated to asymmetric changes in accretion. The mass accretion rates inferred from modeling of L1527 show it to be (Tobin et al., 2010; Flores-Rivera et al., 2021). These values are two orders of magnitude higher when compared to values from Class II protostars, indicating that L1527 is still young. Recent JWST NIRspec observational estimate on mass accretion rate show it to be and supports a non-steady and asymmetric accretion scenario in L1527 (Drechsler et al., 2026).
@@
A key morphological characteristic of L1527 when compared to other Class 0/I protostars is the poor collimation of the outflow and jet emission. It is still unclear whether collimation of jet is a factor of age or environment. However, the common consensus is that as protostars evolve over time, the degree of collimation tends to decrease. This trend is also accompanied by a change in jet properties, such as transitioning from predominantly molecular emission to primarily ionized emission (Frank et al., 2014; Bally , 2016; Mottram et al., 2017). In the case of L1527, we do not see any evidence of a molecular jet in the MIRI/MRS H2 integrated intensity maps (see Fig. 3). Instead, most of the molecular emission is tracing the poorly collimated disk wind. We measured the outflow opening angle using the MIRI broadband images by tracing the cavity walls from about 2000 au down to less than than 50 au near the protostar. The estimated opening angle is about considering no correction for inclination as the outflow is almost in the plane-of-sky. This is consistent with previous estimate of by Tobin et al. (2008) using Spitzer IRAC images. Dunham et al. (2024) studied opening angles of 46 outflows in Perseus cloud and found that outflows widen during the Class 0 phase as they clear dense envelope material, but reach a limit by Class I when most of the envelope has been dispersed and further expansion is no longer supported.
@@
The unusually wide opening angle and lack of a well-collimated molecular jet makes L1527 a unique jet-launching system and suggests either an atypical evolution pathway or the combination of MHD processes and environmental conditions. In particular, magnetocentrifugal winds are thought to self-collimate due to stresses generated by magnetic field lines anchored in the disk or the disk-star interface (Shu et al., 1994; Pudritz et al., 2007). A strong poloidal magnetic field near the disk is responsible for the initial launch and collimation. As the flow propogates, the disk rotation twists the field lines, generating toroidal magnetic field that further collimates the outflow at larger distances (Shang et al., 2006). When the poloidal field near the disk is weak due to turbulence, envelope infall or disk misalignment, the initial outflow will be poorly collimated and subsequent buildup of toroidal fields may not be sufficient to produce strong collimation. Numerical simulations (e.g. Tomisaka , 1998; Hennebelle & Fromang , 2008; Joos et al., 2012) demonstrate that when the poloidal field strength is low or the field is strongly misaligned with the rotation axis, the resulting outflows are wide and poorly collimated.
@@
Segura-Cox et al. (2015) carried out subarcsecond polarization observations of L1527 at 1.3 mm and found a dominant toroidal field in the region without any poloidal component. In addition, they found the magnetic field to be misaligned with the rotation axis. However, recent high sensitivity polarization studies suggest that such polarized emission may primarily trace scattered light and not the intrinsic magnetic field structure (Harrison et al., 2021; Lin et al., 2024). Nevertheless, if taken as a direct tracer, this result combined with our observed morphology is consistent with a scenario of a weak misaligned poloidal field responsible for poor collimation. The strength of the magnetic field in the region also plays an important role in the degree of collimation. Several observations using dust polarization and Zeeman splitting have shown that Class 0/I protostars with highly collimated jets have magnetic field strengths in the order of 1-100 mG at scales less than 100 au, for e.g. IRAS 16293-2422, HH 211 and L1448 IRS2 (Alves et al., 2012; Lee et al., 2018; Kwon et al., 2019). In contrast, in protostars like IRAS 15398-3359 (Redaelli et al., 2019), where the outflows are less collimated, inferred field strengths at similar scales are lower, on the order of 10-100 G. So far there has not been any direct measurement of magnetic field strength in L1527. However, Davidson et al. (2014) tested different magnetic field models with observed polarization configurations to reproduce the observed disk and outflow morphology. They found that only the weakly aligned magnetic field model with strengths of about G favors the observed morphology. Therefore, it is likely that the magnetic field in L1527 is weak and contributes to the poor collimation leading to broad, wide-angle flows rather than narrow jets. These properties in L1527 suggest that even in the earliest embedded stages, variation in magnetic field configuration, in addition to environmental conditions, strongly influence the collimation and shaping of the outflow structures.
@@ 6 Summary and Conclusions
We present JWST MIRI/MRS spectral observations towards a young low-mass Class 0/I protostar: L1527. This study has provided a detailed characterization of the physical and kinematic properties within the innermost 500 au region of the L1527 outflow. Through high-resolution spectroscopic observations we identified several molecular, atomic and ionized lines tracing distinct outflow components. The most striking result observed is the presence of a short and poorly collimated high velocity ionized jet in the eastern region, previously undetected, which highlights the accretion onto the protostar. Conversely, no molecular H2 jet is detected in the MIRI data at smaller scales, nor a CO jet at larger scales in other studies, making L1527 outflow a unique flow among flows from young stars. Below, we summarize the main results of our study.
- Broadband MIR imaging of L1527 reveals a bright bipolar structure consisting of several filaments, bow shocks, knots and reverse shocks, dominated by scattered light. The outflow lobes are significantly bright and have a large opening angle of .
- The MIRI/MRS spectra reveal a chemically rich environment composed of molecular, atomic and ionized species. Pure rotational H2 (0-0) transitions from S(1) to S(8) are observed, with S(4) and S(5) being the strongest. The lines appear to be associated to a single component of the outflow, most likely the outflow cavity wall.
@@
- [Ne ii] channel maps highlight spatially distinct emission regions, indicating shock activity. The radial velocity map shows the eastern lobe is redshifted and the western lobe blueshifted, contrary to earlier interpretations. The average radial velocity was found to be km s-1.
- Double peaked emission profiles in [Ar ii], [Ne iii], and [Fe ii] lines were detected and found confined only to the eastern region. The LV (30 km s-1) component traces the material associated with the general disk wind, while the HV (100 km s-1) component traces a poorly collimated fast jet extending upto 500 au from the protostar.
- Excitation diagram analysis of H2 rotational lines reveal two distinct gas layers: a warm component (550 K, cm-2) and a hot component (2500 K, cm-2). The spatial variations in temperature and column density maps closely follow the outflow morphology in H2 emission, suggesting that gas heating and compression are strongly linked to a common excitation mechanism.
- Comparison of molecular and ionized emission with shock models indicate that moderate velocity -shocks with modest UV field are the dominant driver for the widespread excitation of the observed lines.
Overall, the JWST MIRI/MRS observations of L1527 offers new insights into the early phases of protostellar evolution. The data reveal a chemically rich environment shaped by dynamic interactions and shocks. The outflow exhibits a stratified structure of molecular and ionized gas, where compact jet coexists with a slower wide-angle disk wind, that are likely driven by weak magnetic fields. These findings suggest that stratified wide angle outflows are active even in deeply embedded stages, challenging assumptions that such features are confined to later stages. Future high-spatial-resolution observations combined with magnetic field studies will be essential to further constrain the origin and dynamics of the outflow/jet structures.

### v2

@@ 5 Discussion
JWST observations reveal a detailed hourglass-shaped bipolar outflow in the MIR broadband images (see Fig. 1). The images exhibit multiple nebulous features consisting of warm and hot layers of gas and dust generated within the outflow cavity. Our spectral maps reveal that the emission includes molecular, ionized, and scattered light. Despite such detailed composition, we detect no trace of a large-scale collimated jet in the broadband images. This absence of a large jet signature is distinctive for L1527 given its evolutionary stage. In contrast, recent JWST observations of several other class 0/I protostars, such as HH 211, HH 46, TMC 1, and IRAS 16253-2429, have shown prominent jet emission. Their bulk of the emission is particularly traced in high-J H2 and [Fe ii] lines (Ray et al., 2023; Nisini et al., 2024; Tychoniec et al., 2024; Narang et al., 2024). We see that most of L1527’s dominant emission comes from the wide-angle outflow across both molecular and ionized emission. However, the detection of double-peaked ionized lines with the HV component at velocities 100 km s-1provides compelling evidence for the presence of a jet. This jet signature spans less than 500 au from the protostar and is fainter than the jets from other sources. In addition, compared to the large-scale outflow structures (3000 au), the jet is small and covers only a fraction of the region, thereby concealing its visibility in the broadband images.
@@
Double-peaked emission lines with kinematics similar to L1527 have been observed in other protostellar outflows. For instance, HH 34 presents a collimated jet showing two velocity components both in molecular and ionized lines (Nisini et al., 2016). Using VLT observations of H2 and [Fe ii] lines, Garcia Lopez et al. (2008) showed that the HH 34 LV component is distributed at 15-50 km s-1surrounding a fast axial jet (90-140 km s-1). Likewise, in DG Tau, [Fe ii] and [Ne iii] double-peaked lines were observed with VLT, using SINFONI and X-shooter showing a similar velocity separation (Agra-Amboage et al., 2011; Liu et al., 2016). Many other sources such as SVS 13, HH 211, and HH 46 (Takami et al., 2006; Tabone et al., 2017; Garcia Lopez et al., 2010) exhibit the same pattern. Such profiles are rarely seen in class 0 protostars due to their dense envelopes but are more notable in class II protostars due their moderate extinction. A unifying scenario in all these observations is that the LV component is associated with a slower, wider wind, while the HV component traces the faster jet, thereby pointing to a stratified structure in the outflow.
@@
The observed double-peaked lines can also be interpreted within the framework of shock models. The coexistence of the LV and HV components in the ionized lines suggests a stratified velocity structure within the predominantly -type shock environment. Our comparison of models shows that moderate-velocity ( km s-1) -type shocks illuminated by nonzero UV field best reproduce the observed emission. This suggests that the bulk of line excitation is driven by compact, dissociative shock fronts likely associated with the disk wind. The high [Ne ii]/[Ne iii] across most of the emitting region indicates soft ionization overall, consistent with moderate-velocity shocks where UV photons generated in the shock front contribute to ionization. However, the clear detection of the HV component in [Ar ii], [Ne iii] , and [Fe ii] specifically indicates a localized contribution from a faster jet. The HV component, particularly in [Ne iii] is significant, as neon ionization requires energies 40 eV, typically achieved only via fast dissociative -type shocks (80 km s-1) (Hollenbach & McKee , 1989; Shang et al., 2010), as strong X-ray/UV photoionization is unlikely in L1527. The inferred jet velocity (100 km s-1) in L1527 is consistent with predictions of a shock-excited jet emerging from the inner disk, suggesting the existence of a HV jet core. Together, these observations support a stratified jet-outflow structure in which the fast jet core produces HV component in the ionized lines, while a moderate-velocity -type shock with intermediate UV field contributes to the broader LV component throughout the outflow.
@@
Radio observations have been used to detect jets based on the continuum emission arising from the thermal Bremsstrahlung radiation (free-free emission) (e.g., Anglada et al., 1998; Furuya et al., 2003). An earlier study of L1527 by Reipurth et al. (2004) using the VLA ”A” configuration at 3.6 cm observed an extended continuum jet in the eastern region close to the protostar. The jet emission was extended to about 150 au and was seen to curve toward the south at its leading edge (see the black contours in Fig. 11). This detection was not taken into consideration, because it was suggested that it may be due to a binary, as proposed by Loinard et al. (2002). However, based on recent high-sensitivity observations, it was shown that L1527 is not a binary (e.g., Nakatani et al., 2020; Sheehan et al., 2022). Correlating the morphology of the radio jet to the MIRI jet, we find similarity in the distribution. The jet axis matches spatially, extending toward the east and bending or curving toward the south. The curving of the jet could be due to either wiggling and precession, or an interaction with a dense layer of gas. A key feature that can be interpreted from both observations is the proper motion of the jet. The radio jet extends approximately to about 150 au, whereas the MIRI jet emission extends up to 500 au. There is a difference of about 350 au in length between both the emissions, which have a time difference of 20 years. Based on this, a rough estimate of the average tangential velocity of the jet would lead to around km s-1. A main caveat in this estimate is that both datasets are observed at different wavelengths; therefore, the difference in the jet length may also be associated with the emission from different layers of gas.
@@
More recently, van’t Hoff et al. (2023) carried out ALMA observations and detected a noticeable SiO emission in the western side (see the blue contours in Fig. 11). SiO predominantly traces high-density ejection material associated with jets in the innermost regions of class 0/I protostars (e.g., HH 212; Lee , 2020). The SiO emission observed in L1527 is compact and positioned very close to the protostar (10-50 au). Most importantly, it is blueshifted and exhibits higher velocity when compared to ALMA CO observations. When considered alongside our MIRI jet emission in the eastern side, this detection strongly supports the bipolar nature of the jet in L1527. In addition, their velocities and positions confirm that the eastern region is redshifted while the western region is blueshifted consistent with the [Ne ii] velocity maps (see Fig. 6). van’t Hoff et al. (2023) also examined a large-scale 12CO map and found hints of a collimated CO emission in the western side at a distance of around 500-800 au. This emission appeared more resolved, flowing along the outflow axis, but had lower velocity than the SiO emission. Therefore, it could not be confirmed as a large-scale jet.
@@
While the L1527 outflow exhibits a clear bipolar structure, the jet emission traced by the HV component and SiO appears asymmetric and poorly distributed. The confinement of the HV component to one region may reflect the effects due to combination of inclination, asymmetries in the medium, and/or an episodic accretion. Asymmetric jets have been observed in other deeply embedded sources, such as L1448-C (Dionatos et al., 2009), BHR 71 (Gusdorf et al., 2015), HH 46/47 (Zhang et al., 2016), and TMC 1 (Tychoniec et al., 2024), where external density gradients or variable accretion lead to predominantly one-sided jet activity. The inhomogeneity of the medium significantly changes the excitation conditions leading to emission in only certain tracers for each region. Many theoretical simulations lend strong support to this interpretation. For instance, Raga et al. (2002) modeled jets in young stars and showed how early-stage jets can produce compact, asymmetric jet features, particularly when interacting with clumpy or inhomogeneous envelopes. Hansen et al. (2016) conducted 3D MHD simulations of class 0 jets and found that early jet interactions with over-dense clumps can produce complex asymmetric shock structures. In addition, such interactions also produce a number of reverse shocks along the flow, which seems to be the case in L1527 (see Fig. 1). Moreover, it is believed that magnetic fields play a crucial role in shaping the jet structures (Cerqueira & Gouveia Dal Pino , 2001; Pudritz & Ray , 2019). Variations in magnetic field topology can lead to jets with uneven or asymmetric outflows. Different configurations of magnetic fields, such as misalignments between the field and the rotation axis due to a warped disk, or nonuniform field structures, can cause one side of the jet to be less collimated or more powerful than the other.
@@
Another factor attributing to L1527’s jet morphology is its evolutionary stage. While L1527 is classified as a class 0/I transitional object, its jet appears unusually weak and compact. This indicates that jet activity has either declined or undergoes episodes of periodic mass loss. Such temporal evolution has been seen in sources such as L1448-C, IRAS 16293-2422, and Serpens South C7, where compact jets coexist with broader molecular flows and signs of episodic accretion are common (Bachiller et al., 1990; Kristensen et al., 2011; Plunkett et al., 2015). Spitzer IRAC observations of L1527 taken 12 months apart show total flux variability and emission variability in each outflow lobe (Tobin et al., 2008). Such variations are typically associated with asymmetric changes in accretion. The mass accretion rates inferred from modeling L1527 show it to be (Tobin et al., 2010; Flores-Rivera et al., 2021). These values are two orders of magnitude higher when compared to values from class II protostars, indicating that L1527 is still young. Recent JWST NIRspec observational estimates on mass accretion rate show it to be and support a nonsteady and asymmetric accretion scenario in L1527 (Drechsler et al., 2026).
@@
A key morphological characteristic of L1527, compared to other class 0/I protostars, is the poor collimation of its outflow and jet emission. It is still unclear whether jet collimation is a factor of age or environment. However, the common consensus is that as protostars evolve over time, the degree of collimation tends to decrease. This trend is also accompanied by changes in jet properties, such as the transition from predominantly molecular to primarily ionized emission (Frank et al., 2014; Bally , 2016; Mottram et al., 2017). In the case of L1527, we do not see any evidence of a molecular jet in the MIRI/MRS H2-integrated intensity maps (see Fig. 3). Instead, most of the molecular emission traces the poorly collimated disk wind. We measured the outflow opening angle using the MIRI broadband images by tracing the cavity walls from about 2000 au down to less than than 50 au near the protostar. The estimated opening angle is about , with no inclination correction as the outflow lies nearly in the plane of sky. This is consistent with the previous estimate of by Tobin et al. (2008) using Spitzer IRAC images. Dunham et al. (2024) studied the opening angles of 46 outflows in the Perseus cloud and found that outflows widen during the class 0 phase as they clear dense envelope material, but reach a limit by class I when most of the envelope has been dispersed and further expansion is no longer supported.
@@
The unusually wide opening angle and lack of a well-collimated molecular jet make L1527 a unique jet-launching system and suggest either an atypical evolution pathway or a combination of MHD processes and environmental conditions. In particular, magnetocentrifugal winds are thought to self-collimate due to stresses generated by magnetic field lines anchored in the disk or the disk-star interface (Shu et al., 1994; Pudritz et al., 2007). A strong poloidal magnetic field near the disk is responsible for the initial launch and collimation. As the flow propagates, disk rotation twists the field lines, generating a toroidal magnetic field that further collimates the outflow at larger distances (Shang et al., 2006). When the poloidal field near the disk is weak due to turbulence, envelope infall, or disk misalignment, the initial outflow is poorly collimated, and the subsequent buildup of toroidal fields may not be sufficient to produce strong collimation. Numerical simulations (e.g., Tomisaka , 1998; Hennebelle & Fromang , 2008; Joos et al., 2012) demonstrate that when the poloidal field strength is low or the field strongly misaligns with the rotation axis, the resulting outflows are wide and poorly collimated.
@@
Segura-Cox et al. (2015) carried out subarcsecond polarization observations of L1527 at 1.3 mm and found a dominant toroidal field in the region without any poloidal component. In addition, they found the magnetic field to be misaligned with the rotation axis. However, recent high-sensitivity polarization studies suggest that such polarized emission may primarily trace scattered light rather than the intrinsic magnetic field structure (Harrison et al., 2021; Lin et al., 2024). Nevertheless, if taken as a direct tracer, this result combined with our observed morphology is consistent with a scenario of a weak misaligned poloidal field responsible for poor collimation. The strength of the magnetic field in the region also plays an important role in the degree of collimation. Several observations using dust polarization and Zeeman splitting have shown that class 0/I protostars with highly collimated jets have magnetic field strengths on the order of 1-100 mG at scales less than 100 au, namely IRAS 16293-2422, HH 211, and L1448 IRS2 (Alves et al., 2012; Lee et al., 2018; Kwon et al., 2019). In contrast, in protostars such as IRAS 15398-3359 (Redaelli et al., 2019), where the outflows are less collimated, inferred field strengths at similar scales are lower, on the order of 10-100 G. So far there has not been any direct measurement of magnetic field strength in L1527. However, Davidson et al. (2014) tested different magnetic field models with observed polarization configurations to reproduce the observed disk and outflow morphology. They found that only the weakly aligned magnetic field model with strengths of about G favors the observed morphology. Therefore, it is likely that the magnetic field in L1527 is weak and contributes to poor collimation, leading to broad, wide-angle flows rather than narrow jets. These properties in L1527 suggest that even in the earliest embedded stages, variations in magnetic field configuration, in addition to environmental conditions, strongly influence the collimation and shaping of the outflow structures.
@@ 6 Summary and conclusions
We presented JWST MIRI/MRS spectral observations toward a young low-mass class 0/I protostar: L1527. This study has provided a detailed characterization of the physical and kinematic properties within the innermost 500 au region of the L1527 outflow. Through high-resolution spectroscopic observations, we identified several molecular, atomic, and ionized lines tracing distinct outflow components. The most striking result is a previously undetected, short, and poorly collimated HV ionized jet in the eastern region, which indicates active accretion onto the protostar. Conversely, no molecular H2 jet is detected at smaller scales in the MIRI data, nor a CO jet at larger scales in other studies, making the L1527 outflow a unique among outflows from young stars. Below, we summarize the main results of our study.
- Broadband MIR imaging of L1527 reveals a bright bipolar structure consisting of several filaments, bow shocks, knots, and reverse shocks, dominated by scattered light. The outflow lobes are significantly bright and have a large opening angle of .
- The MIRI/MRS spectra reveal a chemically rich environment composed of molecular, atomic, and ionized species. Pure rotational H2 (0-0) transitions from S(1) to S(8) are observed, with S(4) and S(5) being the strongest. The lines appear associated with a single component of the outflow, most likely the outflow cavity wall.
@@
- [Ne ii] channel maps highlight spatially distinct emission regions, indicating shock activity. The radial velocity map shows that the eastern lobe is redshifted and the western lobe is blueshifted, contrary to earlier interpretations. The average radial velocity is km s-1.
- Double-peaked emission profiles in the [Ar ii], [Ne iii], and [Fe ii] lines were detected and found confined only to the eastern region. The LV (30 km s-1) component traces the material associated with the general disk wind, while the HV (100 km s-1) component traces a poorly collimated fast jet extending up to 500 au from the protostar.
- An excitation diagram analysis of H2 rotational lines reveals two distinct gas layers: a warm component (550 K, cm-2) and a hot component (2500 K, cm-2). The spatial variations in temperature and column density maps closely follow the outflow morphology in H2 emission, suggesting that gas heating and compression are strongly linked to a common excitation mechanism.
- A comparison of molecular and ionized emission with shock models indicates that moderate-velocity -shocks with a modest UV field are the dominant drivers of the widespread excitation seen in the observed lines.
Overall, the JWST MIRI/MRS observations of L1527 offer new insights into the early phases of protostellar evolution. The data reveal a chemically rich environment shaped by dynamic interactions and shocks. The outflow exhibits a stratified structure of molecular and ionized gas, where a compact jet coexists with a slower wide-angle disk wind, likely driven by weak magnetic fields. These findings suggest that stratified wide-angle outflows are active even in deeply embedded stages, challenging assumptions that such features are confined to later stages. Future high-spatial-resolution observations combined with magnetic field studies will be essential to further constrain the origin and dynamics of the outflow and jet structures.